# Breaking the Modality Barrier: Universal Embedding Learning with Multimodal LLMs

### v3

@@ Abstract.
Figure 1. The UniME framework incorporates textual discriminative knowledge distillation and hard negative enhanced instruction tuning stages to learn discriminative representations for diverse downstream tasks. Our framework achieves state-of-the-art performance on both the MMEB benchmark and multiple retrieval tasks.
[…]
[¡short description¿]¡long description¿
@@ 2. Related work
Figure 2. The framework of the Textual Discriminative Knowledge Distillation stage. We leverage the state-of-the-art LLM-based embedding model to enhance the discriminative capabilities of the MLLM’s language component.
[…]
[¡short description¿]¡long description¿
@@ 2.1. Multimodal Large Language Models
$2.$ LLMs for Representation Learning
@@ 3. Method
Figure 3. The framework of the Hard Negative Enhanced Instruction Tuning stage. We further improve the discriminative capabilities of the MLLM through false negative filtering and hard negative sampling.
[…]
[¡short description¿]¡long description¿
This section first establishes the preliminary definitions, including task formulation and feature extraction (Section 3.1). Then we start to introduce our proposed novel two-stage framework UniME. We elaborate on the first textual discriminative knowledge distillation stage in Section 3.2. After that, we present the second hard negative enhanced instruction tuning stage in Section $3.$ Finally, we introduce the training recipe in Section 3.4.
@@ 3.2. Textual Discriminative Knowledge Distillation
$3.$ Hard Negative Enhanced Instruction Tuning
@@ 3.3.2. Hard Negative Sampling
$q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ Replace Swap Add
@@ 4.3.1. Multi-Modal Retrieval
Figure 4. The discrimination comparison between E5-V and UniME. represents the UniME model only training on the first textual discrimination knowledge distillation stage.
[…]
[¡short description¿]¡long description¿
@@ 5.1. Analysis of the Hard Negative
Figure 5. The comparison of training loss and pre-clip gradient norms for hard negatives, easy negatives, and random sample negatives.
[…]
[¡short description¿]¡long description¿
Figure 6. Visualization of the top-k next predicted tokens before and after different training stages based on Phi3.5-V. : UniME with textual discrimination distillation only. : UniME with both textual discrimination distillation and hard negative enhanced instruction tuning.
[…]
[¡short description¿]¡long description¿
@@ 5.4. Ablation on the Number of Hard Negatives
$5.$ Visualization of the Output Distribution
@@ Appendix B External Results
Figure 7. Visualization of the top-k next predicted tokens before and after different training stages based on Phi3.5-V. : UniME with textual discrimination distillation only. : UniME with both textual discrimination distillation and hard negative enhanced instruction tuning.
[…]
[¡short description¿]¡long description¿
@@ B.2. Specific results on the MMEB
Figure 8. Schematic illustration of Negative data.
[…]
[¡short description¿]¡long description¿

### v4

@@ Abstract.
Figure 1. The UniME framework incorporates textual discriminative knowledge distillation and hard negative enhanced instruction tuning stages to learn discriminative representations for diverse downstream tasks. Our framework achieves state-of-the-art performance on both the MMEB benchmark and multiple retrieval tasks.: ¡short description¿
@@ 2. Related work
Figure 2. The framework of the Textual Discriminative Knowledge Distillation stage. We leverage the state-of-the-art LLM-based embedding model to enhance the discriminative capabilities of the MLLM’s language component.: ¡short description¿
@@ 2.1. Multimodal Large Language Models
2.2. LLMs for Representation Learning
@@ 3. Method
Figure 3. The framework of the Hard Negative Enhanced Instruction Tuning stage. We further improve the discriminative capabilities of the MLLM through false negative filtering and hard negative sampling.: ¡short description¿
This section first establishes the preliminary definitions, including task formulation and feature extraction (Section 3.1). Then we start to introduce our proposed novel two-stage framework UniME. We elaborate on the first textual discriminative knowledge distillation stage in Section 3.2. After that, we present the second hard negative enhanced instruction tuning stage in Section 3.3. Finally, we introduce the training recipe in Section 3.4.
@@ 3.2. Textual Discriminative Knowledge Distillation
3.3. Hard Negative Enhanced Instruction Tuning
@@ 3.3.2. Hard Negative Sampling
$q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ Replace Swap Add
@@ 4.3.1. Multi-Modal Retrieval
Figure 4. The discrimination comparison between E5-V and UniME. represents the UniME model only training on the first textual discrimination knowledge distillation stage.: ¡short description¿
@@ 5.1. Analysis of the Hard Negative
Figure 5. The comparison of training loss and pre-clip gradient norms for hard negatives, easy negatives, and random sample negatives.: ¡short description¿
Figure 6. Visualization of the top-k next predicted tokens before and after different training stages based on Phi3.5-V. : UniME with textual discrimination distillation only. : UniME with both textual discrimination distillation and hard negative enhanced instruction tuning.: ¡short description¿
@@ 5.4. Ablation on the Number of Hard Negatives
5.5. Visualization of the Output Distribution
@@ Appendix B External Results
Figure 7. Visualization of the top-k next predicted tokens before and after different training stages based on Phi3.5-V. : UniME with textual discrimination distillation only. : UniME with both textual discrimination distillation and hard negative enhanced instruction tuning.: ¡short description¿
@@ B.2. Specific results on the MMEB
Figure 8. Schematic illustration of Negative data.: ¡short description¿